The SRM oplevs were found to be oscillating because of a small phase margin.
I reduced the gains to the half of them. The peak which Steve found today maybe due to this oscillation.
The SRM bounce peak 18.33 Hz. Suresh helped me to retune filter through Foton, but we failed to remove it.